OpenStudy (anonymous):

x-14 Square root 14+48=0

OpenStudy (anonymous):

you want to find x? ^_^

OpenStudy (anonymous):

im typed incorrect its x-14 square root x+48=0

OpenStudy (anonymous):

\[x-14 = -\sqrt(14+48)\]\[x = -\sqrt(62) + 14\] ^_^

OpenStudy (anonymous):

oh, okay then :)

OpenStudy (anonymous):

\[x - 14 = - \sqrt(x+48)\]\[(x-14)^2 = (x+48)\]\[x-28x+196 = x + 48\]\[x^2-28x - 48 - x +196=0\]\[x^2 - 29x + 148 = 0\] hmmm, there's something wrong

OpenStudy (anonymous):

oh oh wait lol! you can't move sqrt(x+48) to the right since it's not being added or subtracted with (x-14) >_< sorry!

OpenStudy (anonymous):

my bad, sorry ^_^"

OpenStudy (anonymous):

looks like robto is writing down the answer :)

OpenStudy (anonymous):

no sstarica he left

OpenStudy (anonymous):

\[x=-48+14 \sqrt{14} \] if \[48-14 \sqrt{14}+x==0 \]

OpenStudy (anonymous):

the answer has to be in a form like 17,34

OpenStudy (anonymous):

it has to be solutions

OpenStudy (anonymous):

\[48-14\sqrt{14+x}= 0, x=-\frac{110}{49}\]